1. Break-even analysis

To be profitable, a firm must recover its costs. These costs include both its fixed and its variable costs. One way that a firm evaluates at what stage it would recover the invested costs is to calculate how many units or how much in dollar sales is necessary for the firm to earn a profit.

Consider the case of Free Spirit Industries Inc.:

Free Spirit Industries Inc. is considering a project that will have fixed costs of $10,000,000. The product will be sold for $32.50 per unit, and will incur a variable cost of $11.25 per unit.

Given Free Spirit’s cost structure, it will have to sell (470,588 /299,712 / 346,154 / 92,219) units to break even on this project (QBEQBE).

Free Spirit Industries Inc.’s marketing sales director doesn’t think that the market for the firm’s goods is big enough to sell enough units to make the company’s target operating profit of $15,000,000. In fact, she believes that the firm will be able to sell only about 175,000 units. However, she also thinks the demand for Free Spirit Industries Inc.’s product is relatively inelastic, so the firm can increase the sale price. Assuming that the firm can sell 175,000 units, what price must it set to meet the CFO’s EBIT goal of $15,000,000?

$154.11 per unit

$161.82 per unit

$177.23 per unit

$192.64 per unit

What affects the firm’s operating break-even point?

Several factors affect a firm’s operating break-even point. Based on the scenarios described in the following table, indicate whether these factors would increase, decrease, or leave unchanged a firm’s break-even quantity—assuming that only the listed factor changes and all other relevant factors remain constant.

Increase

Decrease

No Change

The firm’s tax rate increases.
The firm depreciates its fixed assets more quickly over a shorter life.
The variable cost per unit decreases.

When a large percentage of a firm’s costs are fixed, the firm is said to have a (Low/High)  degree of operating leverage.
Answer #1

1.
=10000000/(32.5-11.25)=470588.235294118

2.
=(15000000+10000000)/175000+11.25=154.107142857143

3.
No Change

4.
Increase

5.
Increase

6.
High